I need to find the 100th term, for the sequence 2,6,12,20
Answer by Edwin McCravy(20086)



  1st term = 1×2 = 2
  2nd term = 2×3 = 6
  3rd term = 3×4 = 12
  4th term = 4×5 = 20

So we can tell the pattern.  It's (the number of term) times (one more than
the number of term).

Therefore the hundredth term is (the number of term, 100) times (one more than 
the number of term, 101), so

100th term = 100×101 = 10100

In general, the 

Nth term = N×(N+1)

Edwin